# Apache Point rapid response characterization of primitive pre-impact detection asteroid 2024 RW1

### v1

@@ Abstract
Pre-impact detection asteroids (PIDAs) may be detected only a few hours before their impact with Earth, providing a brief opportunity to characterize them before impact. We describe the characterization of PIDA 2024 RW1, which was discovered by the Catalina Sky Survey on 2024 September 4 at 05:43 UTC, before it entered the atmosphere near the northern Philippines at 16:39 UTC. We observed 2024 RW1 with the Astrophysical Research Consortium Telescope Imaging Camera on the Apache Point Astrophysical Research Consortium’s 3.5-m telescope on 2024 September 4 10:16 UTC. We obtained g, r, i, and z photometry of 2024 RW1, yielding color indices of g-r = 0.470.04, r-i = 0.130.04, i-z = -0.110.07, and g-i = 0.600.04, corresponding to a spectral slope of 0.670.40 /100 nm. The closest match to an asteroid spectral type is with B-type asteroids from the C-complex. We detect variations in the time series photometry of the asteroid with an amplitude of 0.75, and a double-peaked rotation period of 1900 s. Assuming a visible albedo of 0.07 and a density of 1500 kg/m3, and using the derived absolute magnitude of 32.20.5, we calculate that the asteroid has a diameter of 1.80.4 m and a total mass of 5000 kg. The most likely source of 2024 RW1 is the 3:1 mean motion resonance followed by the resonance, according to NEOMOD3.
@@ 1 Introduction
Pre-impact detection asteroids (PIDAs) provide an opportunity to test the link between meteorites and their parent bodies in the Main Belt (MB) where they likely originated (Granvik et al., 2016; Chow & Brown, 2025). Characterization of PIDAs using reflected light can be directly compared with fragments that reach the ground as for 2008 TC3 (Jenniskens et al., 2009). The goal of studying PIDAs is to obtain their spectra while still in space, to collect fragments that reach the ground, and contrast high-resolution lab spectra of the meteorites with the parent body. As of 2025 May, only 11 PIDAs have been discovered111https://cneos.jpl.nasa.gov/pi/, accessed 2025 May 14. compared to the 38,000 known near-Earth objects (NEOs) 222https://cneos.jpl.nasa.gov/stats/totals.html, accessed 2025 May 14.. There has been recent success with the recovery of fragments from impacting asteroid 2024 BX1 and the study of its meteor trail (Bischoff et al., 2024; Spurný et al., 2024), but obtaining the combination of pre-atmospheric entry characterization and subsequent collection and analysis of meteor fragments remains challenging.
The majority of PIDAs have been found by all-sky asteroid surveys such as the Catalina Sky Survey (CSS, Larson et al., 1998) and the (Asteroid Terrestrial-impact Last Alert System, ATLAS, Tonry et al., 2018). While surveys provide information on a PIDA’s orbit, dedicated follow-up is required to characterize the asteroids before impact. Most of the PIDAs were discovered less than 24 h before impact (e.g., Jenniskens et al., 2021) which provides a only brief window of time to coordinate multi-wavelength observations of these objects. Rapid-response techniques have been developed to address the characterization of small, close-approaching asteroids, such as multi-filter streak photometry (Bolin et al., 2024) and fast-turnaround triggering of spectroscopic observations at 8-m and 10-m class observatories (Bolin et al., 2025b). Meteor spectrographs also provide an invaluable tool for the characterization of PIDAs as they pass through Earth’s atmosphere (Mazur et al., 2024).
The asteroid 2024 RW1 was discovered by the CSS on 2024 September 04 (Wierzchos et al., 2024), less than 12 h before its impact near Luzon island in the Philippines on 2024 September 04 16:39 (Veres & Green, 2024). This paper describes its observation and characterization using the Astrophysical Research Consortium (ARC) 3.5-m Telescope at Apache Point Observatory (APO). The characterization and analysis of 2024 RW1 will use the spectrophotometric approach of Bolin et al. (2021, 2022, 2025c). We also report the object’s astrometry used to refine its orbit and impact location, and provide its likely source regions within the MB using the NEOMOD3 (Nesvorný et al., 2023).
@@ 2 Observations
Initially designated as CAQTDL2, 2024 RW1 was first detected by the CSS 1.5 m telescope on Mt. Lemmon (Minor Planet Center, MPC, observatory code G96) on 2024 September 04 05:43 UTC. Subsequent observations from both G96 and the Magdalena Ridge Observatory, Socorro 2.4-m telescope (MPC observatory code H01, Pentland et al., 2006) refined its orbital solution and increased its impact probability to 1, setting off alerts from the Center for Near Earth Object Studies (CNEOS) NEO Scout system and the European Space Agency’s Meerkat system (Veres & Green, 2024). They calculated that 2024 RW1 would impact near the northern Philippines at 18.083657∘N, 123.002342∘ E on 2024 September 16:39 UTC.
Following the alerts from NEO Scout, we triggered our observations using the ARC 3.5-m telescope’s Astrophysical Research Consortium Telescope Imaging Camera (ARCTIC) instrument under observing program JH04 (PI: Ingebretsen) at 2024 September 4 09:58 UTC and ended our observations at 2024 September 04 10:57 UTC. ARCTIC has a 7.85 arcmin x 7.85 arcmin field of view and a 0.228 arcsec/pixel scale when binned in 2 x 2 mode (Huehnerhoff et al., 2016). We conducted photometric time series observations of 2024 RW1 using the ARCTIC imaging camera with the Sloan Digital Sky Survey (SDSS) g, r, i, and z filters with effective wavelengths 477 nm, 623 nm, 763 nm and 913 nm respectively (Fukugita et al., 1996). At the beginning of the observations 2024 RW1 was from Earth (Fig. 1), 1.01 au from the Sun, and had a phase angle of 11.6∘. The seeing in the r-band images ranged from 1.4-1.8 arcsec and the airmass of the observations increased from 2.6 to 5.3 through the run.
@@ 3.1 Astrometry
We measured the astrometry using the Astrometrica software (Raab, 2012) and Gaia data release 2 reference stars (Gaia Collaboration et al., 2016, 2018). The astrometry was submitted to the MPC and appeared in Daily Orbit Update R79 (Williams, 2024). In total, 76 observations of 2024 RW1 were submitted between 2024 September 4 5:43 UTC and 2024 September 4 16:02 UTC which were used to fit a final orbit solution by the Jet Propulsion Laboratory (JPL) Horizons system with semi-major axis, = 2.50710802.62 au, eccentricity, = 0.7067752.03, inclination, = 0.52805301.98 degrees, ascending node, = 162.45746351.19 degrees, argument of perihelion, = 249.62235824.01 degrees, mean anomaly, = 349.188161.47 degrees333https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/?sstr=2024%20RW1&view=OPC. The trajectory and final impact locationare shown in Fig. 1.
@@ 3.2 Photometry and spectral classification
The median brightness of 2024 RW1 in the four filters was g = 19.41 0.03, r = 18.94 0.03, i = 18.15 0.03, z = 18.92 0.07 and its color indices are g-r = 0.470.04, r-i = 0.130.04, and i-z = -0.110.07. Its colors are comparable to the Sun in g, r, and i bands (gsun-rsun = 0.44, rsun-isun = 0.11, but it is bluer than the Sun in i-z (isun-zsun = 0.3, Willmer, 2018). The g-i color and spectra slope are 0.600.04 and 0.67/100 nm.
Defining a∗ as a∗ = (0.89 (g-r)) + (0.45 (r-i)) - 0.57 (Ivezić et al., 2001), we find a∗ = -0.090.06 (Fig. 3). 2024 RW1 has a∗ and i-z that are similar to C-types which have a∗ -0.1 and i-z 0.02 but it is significantly bluer than S-types and V-types which have an a∗ 0.13-0.15 (Fig. 4). 2024 RW1 has a relatively small (in magnitude) i-z = -0.110.07 compared to V-types, which have a large (in magnitude) i-z -0.5 due to the presence of strong silicate absorption features at 1 micron (DeMeo & Carry, 2013).
@@ 3.3 Rotational lightcurve
We searched for periodic variations of the r-band lightcurve of 2024 RW1 presented in Table 1. The amplitude of the variations are 0.75, significantly larger than the 0.5-0.1 uncertainties of the individual data points (Fig. 5). We used a similar approach as Bolin et al. (2025a) in testing the periodicity of the lightcurve by using the Lomb-Scargle (L-S) periodogram (Lomb, 1976) (middle panel of Fig. 5) and obtained a single-peaked lightcurve of 976 s. We take a boot strap approach to estimate the lightcurve period and its uncertainty for the r-band data, similar to Bolin et al. (2020) and Purdum et al. (2021), by removing data points from the time series lightcurve and repeating our L-S periodogram estimation of the lightcurve period 10,000 times, resulting in a central value and 1-sigma uncertainty estimate of 979 59 s. We adopted the approach of Bolin & Lisse (2020) by using phase dispersion minimization (PDM) analysis (Stellingwerf, 1978) to check the L-S results with an independent method, resulting in a local minima of 950 s, comparable with the Lomb-Scargle period estimate. Taken together, the L-S and PDM methods imply that 2024 RW1 has a double-peaked rotation period on the order of 1900 s.
@@ 4 Discussion and Conclusions
We can use the spectral and orbital information of 2024 RW1 to estimate its visible albedo, , and its size and MB origin. The spectrum of 2024 RW1 most closely matched the spectra of B-types, which are carbonaceous and have an average visible albedo of 0.07 (Alí-Lagoa et al., 2013). We can then calculate the object’s diameter, , (Harris & Lagerros, 2002) using the absolute magnitude, , of 32.20.5 provided by JPL HORIZONS3, finding that D = 1.80.4 m. Assuming a density of 1500 kg/m3 typical of B-type asteroids (Hanuš et al., 2017), 2024 RW1 ’s mass was 5000 kg. Lastly, we can estimate a rough shape for 2024 RW1 using its observed 0.75 magnitude light curve amplitude. Assuming a triaxial shape with a:b:c and ac, then (Binzel et al., 1989). Using our observed amplitude of = 0.75, we find a 2, implying that 2024 RW1 is an elongated body similar to other solar system and interstellar objects (Mann et al., 2007; Bolin et al., 2018).
We determined the likely MB source of 2024 RW1 from its , , , and H, and the NEOMOD3 NEO population model (Nesvorný et al., 2023, 2024) and found that it is most likely from the 3:1 mean motion resonance with Jupiter, at a 77 probability, followed by the resonance, at a 13 probability, and the 7:3 mean motion resonance, at a 6 probability. Therefore, it seems that 2024 RW1 is from the border between the inner MB and outer MB where several primitive asteroid families are located (Walsh et al., 2013; Delbo et al., 2017). The possible origin of 2024 RW1 being from primitive inner MB families is similar to the likely origin of B-type asteroid (101955) Bennu (Bottke et al., 2015; Takir et al., 2024) suggesting that 2024 RW1 and (101955) Bennu may share a common origin. Future asteroid (e.g., Masiero et al., 2024) and general-purpose all-sky surveys (e.g., Whidden et al., 2019; Vera C. Rubin Observatory LSST Solar System Science Collaboration et al., 2020) may result in the detection of additional PIDAs providing opportunities for the characterization of the Earth-impactor population.

### v2

# Apache Point rapid response characterization of primitive imminent impactor 2024 RW1

@@ Abstract
Imminent impactors may be detected only a few hours before their impact with Earth, providing a brief opportunity to characterize them before impact. We describe the characterization of imminent impactor 2024 RW1, which was discovered by the Catalina Sky Survey on 2024 September 4 at 05:43 UTC, before it entered the atmosphere near the northern Philippines at 16:39 UTC. We observed 2024 RW1 with the Astrophysical Research Consortium Telescope Imaging Camera on the Apache Point Astrophysical Research Consortium’s 3.5-m telescope on 2024 September 4 10:16 UTC. We obtained g, r, i, and z photometry of 2024 RW1, yielding color indices of g-r = 0.470.04, r-i = 0.130.04, i-z = -0.110.07, and g-i = 0.600.04, corresponding to a spectral slope of 0.670.40 /100 nm. The closest match to an asteroid spectral type is with B-type asteroids from the C-complex. We detect variations in the time series photometry of the asteroid with an amplitude of 0.75, and a double-peaked rotation period of 1900 s. Assuming a visible albedo of 0.070.03, a density of 1500 kg/m3, and a calculated absolute magnitude of 30.920.05, we estimate that the asteroid has a diameter of 3.30.7 m and a total mass of 28,000 kg. Comparing our astrometric orbital solutions to NEOMOD3, the most likely source of 2024 RW1 is the 3:1 main belt mean motion resonance (77% probability) followed by the resonance (13% probability), consistent with its organic B-type nature.
@@ 1 Introduction
Imminent impactors provide an opportunity to test the link between meteorites and their parent bodies in the Main Belt (MB) where they likely originated (Granvik et al., 2016; Chow & Brown, 2025). Characterization of imminent impactors using reflected light can be directly compared with fragments that reach the ground as for 2008 TC3 (Jenniskens et al., 2009). The goal of studying imminent impactors is to obtain their spectra while still in space, to collect fragments that reach the ground, and contrast high-resolution lab spectra of the meteorites with the parent body. As of 2025 May, only 11 imminent impactors have been discovered111https://cneos.jpl.nasa.gov/pi/, accessed 2025 May 14. compared to the 38,000 known near-Earth objects (NEOs) 222https://cneos.jpl.nasa.gov/stats/totals.html, accessed 2025 May 14.. There has been recent success with the recovery of fragments from impacting asteroid 2024 BX1 and the study of its meteor trail (Bischoff et al., 2024; Spurný et al., 2024), but obtaining the combination of pre-atmospheric entry characterization and subsequent collection and analysis of meteor fragments remains challenging.
The majority of imminent impactors have been found by all-sky asteroid surveys such as the Catalina Sky Survey (CSS, Larson et al., 1998) and the (Asteroid Terrestrial-impact Last Alert System, ATLAS, Tonry et al., 2018). While surveys provide information on a imminent impactor’s orbit, dedicated follow-up is required to characterize the asteroids before impact. Most of the imminent impactors were discovered less than 24 h before impact (e.g., Jenniskens et al., 2021; Bischoff et al., 2024; Gianotto et al., 2025) which provides a only brief window of time to coordinate multi-wavelength observations of these objects. Rapid-response techniques have been developed to address the characterization of small, close-approaching asteroids, such as the multi-filter streak photometry technique developed by (Bolin et al., 2024) to characterize small NEOs 2016 GE1, 2016 CG18, and 2016 EV84, and which was later used to study Earth impactors, (e.g., 2022 WJ1, 2024 BX1, Devogèle et al., 2024; Kareta et al., 2024; Bolin et al., 2025)and fast-turnaround triggering of spectroscopic observations at 8-m and 10-m class observatories (Bolin et al., 2025a). Meteor spectrographs also provide an invaluable tool for the characterization of imminent impactors as they pass through Earth’s atmosphere (Spurný et al., 2024).
The imminent impactor 2024 RW1 was discovered by the CSS on 2024 September 04 (Wierzchos et al., 2024), less than 12 h before its impact near Luzon island in the Philippines on 2024 September 04 16:39 (Veres & Green, 2024). This paper describes its observation and characterization using the Astrophysical Research Consortium (ARC) 3.5-m Telescope at Apache Point Observatory (APO). The characterization and analysis of 2024 RW1 will use the spectrophotometric approach of Bolin et al. (2021, 2022, 2025b). We also report the object’s astrometry used to refine its orbit and impact location, and provide its likely source regions within the MB using the NEOMOD3 (Nesvorný et al., 2023).
@@ 2 Observations
Initially designated as CAQTDL2, 2024 RW1 was first detected by the CSS 1.5 m telescope on Mt. Lemmon (Minor Planet Center, MPC, observatory code G96) on 2024 September 04 05:43 UTC. Subsequent observations from both G96 and the Magdalena Ridge Observatory, Socorro 2.4-m telescope (MPC observatory code H01, Pentland et al., 2006) refined its orbital solution and increased its impact probability to 1, setting off alerts from the Center for Near Earth Object Studies (CNEOS) Scout system and the European Space Agency’s Meerkat system (Veres & Green, 2024). They calculated that 2024 RW1 would impact near the northern Philippines at 18.083657∘N, 123.002342∘ E on 2024 September 16:39 UTC.
Following the alerts from NEO Scout, we triggered our observations using the ARC 3.5-m telescope’s Astrophysical Research Consortium Telescope Imaging Camera (ARCTIC) instrument under observing program JH04 (PI: Ingebretsen) at 2024 September 4 09:58 UTC and ended our observations at 2024 September 04 10:57 UTC. ARCTIC has a 7.85 arcmin x 7.85 arcmin field of view and a 0.228 arcsec/pixel scale when binned in 2 x 2 mode (Huehnerhoff et al., 2016). We conducted photometric time series observations of 2024 RW1 using the ARCTIC imaging camera with the Sloan Digital Sky Survey (SDSS) g, r, i, and z filters with effective wavelengths 477 nm, 623 nm, 763 nm and 913 nm respectively (Fukugita et al., 1996). At the beginning of the observations 2024 RW1 was from Earth (Fig. 1), 1.01 au from the Sun, and had a phase angle of 11.6∘. The seeing in the r-band images ranged from 1.4-1.8 arcsec and the airmass of the observations increased from 2.6 to 5.3 through the ARC 3.5m observing run.
@@ 3.1 Astrometry
We measured the astrometry using the Astrometrica software (Raab, 2012) and Gaia data release 2 reference stars (Gaia Collaboration et al., 2016, 2018). The astrometry was submitted to the MPC and appeared in Daily Orbit Update R79 (Williams, 2024). In total, 76 observations of 2024 RW1 were submitted between 2024 September 4 5:43 UTC and 2024 September 4 16:02 UTC which were used to fit a final orbit solution by the Jet Propulsion Laboratory (JPL) Horizons system. The solution from the date 2024 September 09 09:22333https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/?sstr=2024%20RW1&view=OPC) is shown in Table 1. The trajectory and final impact location are shown in Fig. 1.
| Heliocentric Elements | ||
| Epoch (JD) | 2,460,557.5 | |
| Time of perihelion, (JD) | 2,460,601.046693 | (5.9304x10-4) |
| Semi-major axis, (au) | 2.507108 | (2.039x10-4) |
| Eccentricity, | 0.7067746 | (2.627x10-5) |
| Perihelion, (au) | 0.73514815 | (6.0749x10-6) |
| Aphelion, (au) | 4.279068 | (3.480x10-4) |
| Inclination, (∘) | 0.5280530 | (1.9849x10-5) |
| Ascending node, (∘) | 162.4574635 | (1.1943x10-5) |
| Argument of perihelion, (∘) | 249.6223582 | (4.0094x10-5) |
| Mean Anomaly, (∘) | 349.18816 | (1.466x10-3) |
@@ 3.2 Photometry and spectral classification
The median brightness of 2024 RW1 in the four filters was g = 19.41 0.03, r = 18.94 0.03, i = 18.15 0.03, z = 18.92 0.07 and its color indices are g-r = 0.470.04, r-i = 0.130.04, and i-z = -0.110.07. Its colors are comparable to the Sun in g, r, and i bands (gsun-rsun = 0.44, rsun-isun = 0.11, but it is bluer than the Sun in i-z (isun-zsun = 0.3, Willmer, 2018). The g-i color and spectra slope are 0.600.04 and 0.67/100 nm. All error estimates are 1-sigma.
Defining a∗ as a∗ = (0.89 (g-r)) + (0.45 (r-i)) - 0.57 (Ivezić et al., 2001), we find a∗ = -0.090.06 (Fig. 3). 2024 RW1 has a∗ and i-z that are similar to C-types which have a∗ -0.1 and i-z 0.02 but it is significantly bluer than S-types and V-types which have an a∗ 0.13-0.15 (Fig. 4). 2024 RW1 has a relatively small (in magnitude) i-z = -0.110.07 compared to V-types, which have a large (in magnitude) i-z -0.5 due to the presence of strong silicate absorption features at 1 m (DeMeo & Carry, 2013).
@@
We estimate the absolute magnitude, , using our g and r observations of 2024 RW1. First we convert the our g and r magnitudes to a V magnitude using a transformation equation from Jester et al. (2005),
We calculate a V magnitude of . Then we can calculate the absolute magnitude with the following equation
taken from Bowell et al. (1988). Here, is the heliocentric distance of 1.011 au, is the geocentric distance of 0.003 au, is the phase angle of 10.872∘ that 2024 RW1 had during our observations, is the phase coefficient which we adopt the value of 0.03, the average value for C-type asteroids (Vereš et al., 2015), and and are the basis functions normalized at = 0∘ described in (Bowell et al., 1988). Using the above equation, we estimate that = 30.920.05, though our uncertainty on is underestimated due to the unknown phase function of 2024 RW1.
@@ 3.3 Rotational lightcurve
We searched for periodic variations of the r-band lightcurve of 2024 RW1 presented in Table 2. The amplitude of the variations are 0.75, significantly larger than the 0.5-0.1 uncertainties of the individual data points (Fig. 5). We used a similar approach as Bolin et al. (2025) in testing the periodicity of the lightcurve by using the Lomb-Scargle (L-S) periodogram (Lomb, 1976) (middle panel of Fig. 5) and obtained a single-peaked lightcurve of 976 s. We take a boot strap approach to estimate the lightcurve period and its uncertainty for the r-band data, similar to Bolin et al. (2020) and Purdum et al. (2021), by removing data points from the time series lightcurve and repeating our L-S periodogram estimation of the lightcurve period 10,000 times, resulting in a central value and 1-sigma uncertainty estimate of 979 59 s. We adopted the approach of Bolin & Lisse (2020) by using phase dispersion minimization (PDM) analysis (Stellingwerf, 1978) to check the L-S results with an independent method, resulting in a local minima of 950 s, comparable with the Lomb-Scargle period estimate. Taken together, the L-S and PDM methods imply that 2024 RW1 has a double-peaked rotation period on the order of 1900 s.
. Bottom panel: phase dispersion minimization analysis of lightcurve rotation period vs. metric (Stellingwerf, 1978). The metric is minimized at a single-peaked rotation period of 950.4 s implying a double-peaked rotation period of 1900.8 s.
@@ 4 Discussion and Conclusions
We can use the spectral and orbital information of 2024 RW1 to estimate its visible albedo, , and its size and MB origin. The spectrum of 2024 RW1 most closely matched the spectra of B-type asteroids, which are carbonaceous and have an average visible albedo of 0.070.03 (Al´ı-Lagoa et al., 2013). We can then calculate the object’s diameter, , (Harris & Lagerros, 2002) using the absolute magnitude, , of 30.920.05 from Section 3.2, finding that D = 3.30.7 m. Assuming a density of 1500 kg/m3 typical of B-type asteroids (Hanuš et al., 2017), 2024 RW1 ’s mass was 28,000 kg. Lastly, we can estimate a rough shape for 2024 RW1 using its observed 0.75 magnitude light curve amplitude. Assuming a triaxial shape with a:b:c and ac, then (Binzel et al., 1989). Using our observed amplitude of = 0.75, we find a 2, implying that 2024 RW1 is elongated similar to other minor bodies (Hudson & Ostro, 1999; Mann et al., 2007; Bolin et al., 2018).
We determined the likely MB source of 2024 RW1 from its , , , and H, and the NEOMOD3 NEO population model (Nesvorný et al., 2023, 2024) and found that it is most likely from the 3:1 mean motion resonance with Jupiter, at a 77 probability, followed by the resonance, at a 13 probability, and the 7:3 mean motion resonance, at a 6 probability. Therefore, it seems that 2024 RW1 is from the border between the inner MB and outer MB where several primitive asteroid families are located (Walsh et al., 2013; Delbo et al., 2017). Current and future asteroid (e.g., Larson et al., 1998; Denneau et al., 2013; Tonry et al., 2018; Masiero et al., 2024) and general-purpose all-sky surveys (e.g., Whidden et al., 2019; Vera C. Rubin Observatory LSST Solar System Science Collaboration et al., 2020) may result in the detection of additional imminent impactors providing opportunities for the characterization of the Earth-impactor population.